# Thermal Modelling of Battery Cells for Optimal Tab and Surface Cooling Control

### v1

@@ III-D3 Model decomposition for independent and targeted cooling control
The result obtained in (34)–(35) implies that for the particular problem in (13), which leads to the particular solution $T_{p}$ and features linearity in two spatial dimensions, the solution trajectory in response to the cooling power applied at each individual boundary satisfies the superposition principle. Additionally, the particular solution ${T}_{p}(\tilde{r},\tilde{z},t)$ at each position $(\tilde{r},\tilde{z})$ can be expressed as a linear combination of the cooling power applied at different boundaries.
@@ III-E State-space model representation
Substituting the particular solution in (34)–(35) into (12) and using the linearity of the partial differential operator, the term $q^{*}$, which is required for deriving the homogeneous solution $T_{h}$, can be accordingly obtained
@@
Based on the results in Section III-CIII-D, a control-oriented 2D thermal model of cylindrical battery cells is achieved. We define $X$ as the state variable, $u$ the input, $Y$ the output, and $w$ the disturbance. According to (19), $X$ is given by
@@
The input $u$ is composed of the cooling power applied at the different places, leading to $u=\left[u_{s},\>u_{t},\>u_{b}\right]^{T}$. The disturbance $w$ represents the heat generation rate, namely $w=q$. The outputs of interest $Y$, include the temperatures at the mid-points of the surface, core, top, and bottom sides. Namely, there exists
$Y(t)=\begin{bmatrix}\hat{T}(1,0,t)&\hat{T}(-1,0,t)&\hat{T}(0,1,t)&\hat{T}(0,-1,t)\end{bmatrix}^{T},$ (38)
@@
$G\dot{X}(t)=\>$ $AX(t)+Bu(t)+Fw(t),$ (39a)
$Y(t)=\>$ $CX(t)+\hat{T}_{p}(t),$ (39b)
where with the additional term $\hat{T}_{p}=[\hat{T}_{p}^{s},\>\hat{T}_{p}^{c},\>\hat{T}_{p}^{t},\>\hat{T}_{p}^{b}]^{T}$, the matrices $G$, $A$, $B$, $C$, and the vector $F$ are derived from (9), (12), (19), and (34)–(35)
$G=\>$ $\rho C_{p}\big{\langle}r\phi_{m}^{\tilde{r}}\phi_{n}^{\tilde{z}},\eta\big{\rangle},$ (40a)
$A=\>$ $\bigg{\langle}r\bigg{[}\alpha^{2}k_{r}\frac{\partial^{2}\phi_{m}^{\tilde{r}}\phi_{n}^{\tilde{z}}}{\partial\tilde{r}^{2}}+\gamma\frac{\partial\phi_{m}^{\tilde{r}}\phi_{n}^{\tilde{z}}}{\partial\tilde{r}}$
@@
$B(:,1)=\>$ $\bigg{\langle}r\bigg{[}\alpha^{2}k_{r}\frac{\partial^{2}\hat{T}_{p}^{s}}{\partial\tilde{r}^{2}}+\gamma\frac{\partial\hat{T}_{p}^{s}}{\partial\tilde{r}}+\beta^{2}k_{z}\frac{\partial^{2}\hat{T}_{p}^{s}}{\partial\tilde{z}^{2}}\bigg{]},\eta\bigg{\rangle},$ (40c)
$B(:,2)=\>$ $\bigg{\langle}r\bigg{[}\alpha^{2}k_{r}\frac{\partial^{2}\hat{T}_{p}^{t}}{\partial\tilde{r}^{2}}+\gamma\frac{\partial\hat{T}_{p}^{t}}{\partial\tilde{r}}+\beta^{2}k_{z}\frac{\partial^{2}\hat{T}_{p}^{t}}{\partial\tilde{z}^{2}}\bigg{]},\eta\bigg{\rangle},$ (40d)
$B(:,3)=\>$ $\bigg{\langle}r\bigg{[}\alpha^{2}k_{r}\frac{\partial^{2}\hat{T}_{p}^{b}}{\partial\tilde{r}^{2}}+\gamma\frac{\partial\hat{T}_{p}^{b}}{\partial\tilde{r}}+\beta^{2}k_{z}\frac{\partial^{2}\hat{T}_{p}^{b}}{\partial\tilde{z}^{2}}\bigg{]},\eta\bigg{\rangle},$ (40e)
$C(1,:)=\>$ $\phi_{m}^{\tilde{r}}(1)\phi_{n}^{\tilde{z}}(0),\>\>C(2,:)=\phi_{m}^{\tilde{r}}(-1)\phi_{n}^{\tilde{z}}(0),$ (40f)
$C(3,:)=\>$ $\phi_{m}^{\tilde{r}}(0)\phi_{n}^{\tilde{z}}(1),\>\>C(4,:)=\phi_{m}^{\tilde{r}}(0)\phi_{n}^{\tilde{z}}(-1),$ (40g)
$F=\>$ $\langle r,\eta\rangle.$ (40h)
@@ III-F Application of the modelling techniques to pouch cells
where the scaling factors $\lambda=2/D$ and $\zeta=2/H$. The new spatial coordinates $\tilde{x}=2x/D-1$ and $\tilde{y}=2y/H-1$. The above dynamic equation is subject to the boundary conditions
@@
$u_{fs}(t)=\>$ $h_{fs}(t)T_{fs,\infty}(t),\>\>\>\>\>\>\>u_{bs}(t)=-h_{bs}(t)T_{bs,\infty}(t),$ (43a)
$u_{t}(t)=\>$ $h_{t}(t)T_{t,\infty}(t),\>\>\>\>\>\>\>\>\>\>\>\>u_{b}(t)=-h_{b}(t)T_{b,\infty}(t).$ (43b)
The outputs of interest $Y$, for pouch battery cells are located at the middle of each side, with the mathematical definition the same as (38).
@@
$G$ $=\rho C_{p}\big{\langle}\phi_{m}^{\tilde{x}}\phi_{n}^{\tilde{y}},\eta\big{\rangle},$ (44a)
$A$ $=\bigg{\langle}\lambda^{2}k_{x}\frac{\partial^{2}\phi_{m}^{\tilde{x}}\phi_{n}^{\tilde{y}}}{\partial\tilde{x}^{2}}+\zeta^{2}k_{y}\frac{\partial^{2}\phi_{m}^{\tilde{x}}\phi_{n}^{\tilde{y}}}{\partial\tilde{y}^{2}},\eta\bigg{\rangle},$ (44b)
$B(:\>,1)$ $=\bigg{\langle}\lambda^{2}k_{x}\frac{\partial^{2}\hat{T}^{fs}_{p}}{\partial\tilde{x}^{2}}+\zeta^{2}k_{y}\frac{\partial^{2}\hat{T}^{fs}_{p}}{\partial\tilde{y}^{2}},\eta\bigg{\rangle},$ (44c)
$B(:,2)$ $=\bigg{\langle}\lambda^{2}k_{x}\frac{\partial^{2}\hat{T}^{bs}_{p}}{\partial\tilde{x}^{2}}+\zeta^{2}k_{y}\frac{\partial^{2}\hat{T}^{bs}_{p}}{\partial\tilde{y}^{2}},\eta\bigg{\rangle},$ (44d)
$B(:\>,3)$ $=\bigg{\langle}\lambda^{2}k_{x}\frac{\partial^{2}\hat{T}^{t}_{p}}{\partial\tilde{x}^{2}}+\zeta^{2}k_{y}\frac{\partial^{2}\hat{T}^{t}_{p}}{\partial\tilde{y}^{2}},\eta\bigg{\rangle},$ (44e)
$B(:\>,4)$ $=\bigg{\langle}\lambda^{2}k_{x}\frac{\partial^{2}\hat{T}^{b}_{p}}{\partial\tilde{x}^{2}}+\zeta^{2}k_{y}\frac{\partial^{2}\hat{T}^{b}_{p}}{\partial\tilde{y}^{2}},\eta\bigg{\rangle},$ (44f)
$C(1,:)$ $=\phi_{m}^{\tilde{x}}(1)\phi_{n}^{\tilde{y}}(0),\>\>C(2,:)=\phi_{m}^{\tilde{x}}(-1)\phi_{n}^{\tilde{y}}(0),$ (44g)
$C(3,:)$ $=\phi_{m}^{\tilde{x}}(0)\phi_{n}^{\tilde{y}}(1),\>\>C(4,:)=\phi_{m}^{\tilde{x}}(0)\phi_{n}^{\tilde{y}}(-1),$ (44h)
$F$ $=\langle 1,\eta\big{\rangle},$ (44i)
@@ IV-B Specification and simulation setup
To demonstrate the modelling for cylindrical cells, we choose a large format lithium-iron-phosphate (LFP) $45$Ah cell and take the corresponding model parameters from [25]. Specifically, for dimension-related parameters, we have $L$ = 198 mm, $R_{\text{out}}$ = 32 mm, $R_{\text{in}}$ = 4 mm [25], and for thermo-physical parameters, there exist $\rho$ = 2118 kgm-3, $C_{p}$ = 795 Jkg-1K-1 $k_{r}$ = 0.67 Wm-1K-1, and $k_{z}$ = 66.6 Wm-1K-1 [26, 11]. For the TEC model of this cylindrical cell, $C_{s}$ is found to be 48.35 JK-1. Regarding $R_{c}$ and $R_{u}$, it is important to note that different sets of parameters will be obtained, depending on the employed cooling scenario. For brevity, only the identified parameters for surface cooling are presented here. After validation, we obtained the parameter values as $C_{c}$ = 1079.6 JK-1, $R_{c}$ = 0.65 KW-1, and $R_{u}$ = 0.08 KW-1.
@@
where $I,V,V_{\text{ocv}}$, and $V_{b}$ are the battery current, terminal voltage, open-circuit voltage, and volume, respectively. Here, the trajectories of $I$ and $V$ depend on the battery usage. We select the worldwide harmonised light vehicle test procedure (WLTP) [28] and scale it by a factor of two for heat generation, which is shown in Fig. 2a. It represents a wide range of vehicle driving conditions on urban, suburban, and highway roads. Based on where the cooling action is imposed, five different cooling scenarios are investigated, including surface cooling (SC), bottom tab cooling (bTC), bottom tab and surface cooling (bTSC), bottom and top tabs cooling (btTC), and all-tabs and surface cooling (aTSC). aTSC has all sides of the battery cell exposed to cooling and is representative of immersion cooling [29] where the cell is immersed in a dielectric fluid. Forced convection liquid cooling (for instance, using a cooling plate) via water or glycol is considered to apply on areas where tab or surface cooling occurs. This typically has a large convection coefficient $h$, which is set as 400 Wm-2K-1 according to [4]. For areas not exposed to active cooling, mild air convection occurs, resulting in a small convection coefficient specified as $h$ = 30 Wm-2K-1.
All the simulations are conducted on a MacBook Pro with a 2.6GHz 6-Core Intel Core i7 processor and 16GB RAM. The first type of benchmark model is solved by the finite element method with an extremely fine mesh of 3610 triangular elements and is implemented in COMSOL Multiphysics v6.0 [30]. Our developed CSG models and their TEC alternatives are implemented in MATLAB R2023b. For simplicity in simulating the CSG models, the numbers of basis functions in the radial and axial directions, as introduced in (15), are chosen to be equal. Consequently, there exists $M=N$ and the model order $O=N^{2}$.
@@ IV-C Results of validation against the FEM model
The examination of the developed modelling method and their underlying assumptions for control-oriented applications is conducted for the battery cells and cooling scenarios specified in Section IV-B. Different CSG model orders, including $O$ = 1, 4, 9, 16, and 25, were evaluated. The results for cylindrical cell modelling compared with the ground truth represented by the FEM model are presented in Fig. 2.
Figure 2: Validation of the developed CSG thermal model for cylindrical battery cells. (a) Heat generation profile of the WLTP. (b) Comparison of the CSG model of orders $O$ = 1 and 4, against the FEM model with the element number $N_{e}$ = 3610 under surface cooling (SC). Comparisons are done for the surface and core mid-point temperatures, i.e., $\hat{T}(1,0,t)$ and $\hat{T}(-1,0,t)$. (c) Comparison of the CSG thermal model against FEM under the scenario of all-tabs and surface cooling (aTSC). (d) Maximum error of the CSG model with different orders under all cooling cases.
According to Fig. 2d, the developed CSG model with an order of 25 can perfectly capture the battery cell’s distributed thermal dynamics under any of the considered cooling scenarios. The resulting modelling error is bounded by 0.03 C. Reducing the model order from 25 to 9, the maximum temperature deviation between the model’s output and its ground truth is 0.4 C within the entire test process. Similar observations are obtained for the studied pouch cell, in which the prediction error for the model with an order of 9 is less than 0.6 C. Remarkably, this high level of model fidelity is achieved in the presence of aggressive driving and cooling. The prediction error gradually increases when the model order is further lowered. This is consistent with Remark 2 and the applied finite-sum approximations in (14) and (24). Indeed, the model with an order of 1 or 4 can still well follow the trajectories of the true surface and core temperatures under different cooling scenarios, as seen in Fig. 2b and c. Specifically, for SC, which involves one-side cooling, $O$ = 1 results in a maximum error of 1.26 C. For aTSC, which involves resolution in multiple directions (multi-side cooling), the same order produces a maximum error of 2.10 C. It is worth mentioning that with this magnitude of local errors under all cooling scenarios, the CSG model can satisfy a range of battery management applications.

### v2

@@ III-D3 Model decomposition for independent and targeted cooling control
The result obtained in (33)–(34) implies that for the particular problem in (13), which leads to the particular solution $T_{p}$ featuring linearity in two spatial dimensions, the solution trajectory in response to the cooling power applied at each individual boundary satisfies the superposition principle. Additionally, the particular solution ${T}_{p}(\tilde{r},\tilde{z},t)$ at each position $(\tilde{r},\tilde{z})$ can be expressed as a linear combination of the cooling powers applied at different boundaries.
@@ III-E State-space model representation
Substituting the particular solution in (33)–(34) into (12) and using the linearity of the partial differential operator, the term $q^{*}$, which is required for deriving the homogeneous solution $T_{h}$, can be obtained accordingly
@@
Based on the results in Section III-CIII-D, a control-oriented 2D thermal model of cylindrical battery cells is achieved. We define $X$ as the state variable, $u$ the input, $\mathcal{Y}$ the output, and $w$ the disturbance. According to (19), $X$ is given by
@@
[…]
$\mathcal{Y}(t)=\begin{bmatrix}\hat{T}(1,0,t)&\hat{T}(-1,0,t)&\hat{T}(0,1,t)&\hat{T}(0,-1,t)\end{bmatrix}^{T},$ (37)
@@
$\mathcal{G}\dot{X}(t)=\>$ $\mathcal{A}X(t)+\mathcal{B}u(t)+\mathcal{F}w(t),$ (38a)
$\mathcal{Y}(t)=\>$ $\mathcal{C}X(t)+\mathcal{D}u(t),$ (38b)
where with the additional term $\mathcal{D}=[\hat{T}_{p}^{s},\>\hat{T}_{p}^{c},\>\hat{T}_{p}^{t},\>\hat{T}_{p}^{b}]^{T}$, the matrices $\mathcal{G}$, $\mathcal{A}$, $\mathcal{B}$, $\mathcal{C}$, and the vector $\mathcal{F}$ are derived from (9), (12), (19), and (33)–(34) to yield
$\mathcal{G}=\>$ $\rho c_{p}\big{\langle}r\phi_{m}^{\tilde{r}}\phi_{n}^{\tilde{z}},\eta\big{\rangle},$ (39a)
$\mathcal{A}=\>$ $\bigg{\langle}r\bigg{[}\alpha^{2}k_{r}\frac{\partial^{2}\phi_{m}^{\tilde{r}}\phi_{n}^{\tilde{z}}}{\partial\tilde{r}^{2}}+\gamma\frac{\partial\phi_{m}^{\tilde{r}}\phi_{n}^{\tilde{z}}}{\partial\tilde{r}}$
@@
$\mathcal{B}(:,1)=\>$ $\bigg{\langle}r\bigg{[}\alpha^{2}k_{r}\frac{\partial^{2}\hat{T}_{p}^{s}}{\partial\tilde{r}^{2}}+\gamma\frac{\partial\hat{T}_{p}^{s}}{\partial\tilde{r}}+\beta^{2}k_{z}\frac{\partial^{2}\hat{T}_{p}^{s}}{\partial\tilde{z}^{2}}\bigg{]},\eta\bigg{\rangle},$ (39c)
$\mathcal{B}(:,2)=\>$ $\bigg{\langle}r\bigg{[}\alpha^{2}k_{r}\frac{\partial^{2}\hat{T}_{p}^{t}}{\partial\tilde{r}^{2}}+\gamma\frac{\partial\hat{T}_{p}^{t}}{\partial\tilde{r}}+\beta^{2}k_{z}\frac{\partial^{2}\hat{T}_{p}^{t}}{\partial\tilde{z}^{2}}\bigg{]},\eta\bigg{\rangle},$ (39d)
$\mathcal{B}(:,3)=\>$ $\bigg{\langle}r\bigg{[}\alpha^{2}k_{r}\frac{\partial^{2}\hat{T}_{p}^{b}}{\partial\tilde{r}^{2}}+\gamma\frac{\partial\hat{T}_{p}^{b}}{\partial\tilde{r}}+\beta^{2}k_{z}\frac{\partial^{2}\hat{T}_{p}^{b}}{\partial\tilde{z}^{2}}\bigg{]},\eta\bigg{\rangle},$ (39e)
$\mathcal{C}(1,:)=\>$ $\phi_{m}^{\tilde{r}}(1)\phi_{n}^{\tilde{z}}(0),\>\>\mathcal{C}(2,:)=\phi_{m}^{\tilde{r}}(-1)\phi_{n}^{\tilde{z}}(0),$ (39f)
$\mathcal{C}(3,:)=\>$ $\phi_{m}^{\tilde{r}}(0)\phi_{n}^{\tilde{z}}(1),\>\>\mathcal{C}(4,:)=\phi_{m}^{\tilde{r}}(0)\phi_{n}^{\tilde{z}}(-1),$ (39g)
$\mathcal{F}=\>$ $\langle r,\eta\rangle.$ (39h)
@@ III-F Application of the modelling techniques to pouch cells
with scaling factors $\lambda=2/D$ and $\zeta=2/H$. The new spatial coordinates are $\tilde{x}=2x/D-1$ and $\tilde{y}=2y/H-1$. The above dynamic equation is subject to the boundary conditions
@@
$u_{fs}(t)=\>$ $h_{fs}(t)T_{fs,\infty}(t),\>\>u_{bs}(t)=-h_{bs}(t)T_{bs,\infty}(t),$ (42a)
$u_{t}(t)=\>$ $h_{t}(t)T_{t,\infty}(t),\>\>u_{b}(t)=-h_{b}(t)T_{b,\infty}(t).$ (42b)
The outputs of interest $\mathcal{Y}$, for pouch battery cells are located at the middle of each side, with the mathematical definition being the same as (37).
@@
$\mathcal{G}$ $=\rho c_{p}\big{\langle}\phi_{m}^{\tilde{x}}\phi_{n}^{\tilde{y}},\eta\big{\rangle},$ (43a)
$\mathcal{A}$ $=\bigg{\langle}\lambda^{2}k_{x}\frac{\partial^{2}\phi_{m}^{\tilde{x}}\phi_{n}^{\tilde{y}}}{\partial\tilde{x}^{2}}+\zeta^{2}k_{y}\frac{\partial^{2}\phi_{m}^{\tilde{x}}\phi_{n}^{\tilde{y}}}{\partial\tilde{y}^{2}},\eta\bigg{\rangle},$ (43b)
$\mathcal{B}(:\>,1)$ $=\bigg{\langle}\lambda^{2}k_{x}\frac{\partial^{2}\hat{T}^{fs}_{p}}{\partial\tilde{x}^{2}}+\zeta^{2}k_{y}\frac{\partial^{2}\hat{T}^{fs}_{p}}{\partial\tilde{y}^{2}},\eta\bigg{\rangle},$ (43c)
$\mathcal{B}(:,2)$ $=\bigg{\langle}\lambda^{2}k_{x}\frac{\partial^{2}\hat{T}^{bs}_{p}}{\partial\tilde{x}^{2}}+\zeta^{2}k_{y}\frac{\partial^{2}\hat{T}^{bs}_{p}}{\partial\tilde{y}^{2}},\eta\bigg{\rangle},$ (43d)
$\mathcal{B}(:\>,3)$ $=\bigg{\langle}\lambda^{2}k_{x}\frac{\partial^{2}\hat{T}^{t}_{p}}{\partial\tilde{x}^{2}}+\zeta^{2}k_{y}\frac{\partial^{2}\hat{T}^{t}_{p}}{\partial\tilde{y}^{2}},\eta\bigg{\rangle},$ (43e)
$\mathcal{B}(:\>,4)$ $=\bigg{\langle}\lambda^{2}k_{x}\frac{\partial^{2}\hat{T}^{b}_{p}}{\partial\tilde{x}^{2}}+\zeta^{2}k_{y}\frac{\partial^{2}\hat{T}^{b}_{p}}{\partial\tilde{y}^{2}},\eta\bigg{\rangle},$ (43f)
$\mathcal{C}(1,:)$ $=\phi_{m}^{\tilde{x}}(1)\phi_{n}^{\tilde{y}}(0),\>\>\mathcal{C}(2,:)=\phi_{m}^{\tilde{x}}(-1)\phi_{n}^{\tilde{y}}(0),$ (43g)
$\mathcal{C}(3,:)$ $=\phi_{m}^{\tilde{x}}(0)\phi_{n}^{\tilde{y}}(1),\>\>\mathcal{C}(4,:)=\phi_{m}^{\tilde{x}}(0)\phi_{n}^{\tilde{y}}(-1),$ (43h)
$\mathcal{F}$ $=\langle 1,\eta\big{\rangle},$ (43i)
@@
At higher C-rates, localized heat generation at the battery tabs can significantly influence the internal temperature distribution of the cell [26]. Future work will explore modelling approaches that treat the tabs and cell body as thermally coupled regions using a contact resistance framework. This will enable a more spatially resolved thermal analysis and better capture the impact of tab heating on overall cell performance.
Due to the internal structure of prismatic cells, significant thermal gradients can develop in all space dimensions ($x$, $y$, and $z$), requiring a 3D thermal model for better accuracy. However, the underlying solution processes for the prismatic cell is equivalent to that of the cylindrical and pouch cells already discussed.
@@ IV-B Specification and simulation setup
To demonstrate the modelling for cylindrical cells, we choose a large format lithium-iron-phosphate (LFP) $45$Ah cell and take the corresponding model parameters from [29]. Specifically, for dimension-related parameters, we have $L$ = 198 mm, $R_{\text{out}}$ = 32 mm, $R_{\text{in}}$ = 4 mm [29], and for thermo-physical parameters, we use $\rho$ = 2118 kgm-3, $c_{p}$ = 795 Jkg-1K-1 $k_{r}$ = 0.67 Wm-1K-1, and $k_{z}$ = 66.6 Wm-1K-1 [30, 13]. For the TEC model of this cylindrical cell, $C_{s}$ is found to be 48.35 JK-1. Regarding $R_{c}$ and $R_{u}$, it is important to note that different sets of parameters will be obtained, depending on the employed cooling scenario. For brevity, only the identified parameters for surface cooling are presented here. After validation, we obtained the parameter values as $C_{c}$ = 1079.6 JK-1, $R_{c}$ = 0.65 KW-1, and $R_{u}$ = 0.08 KW-1.
@@
where $I,V,V_{\text{ocv}}$, and $V_{b}$ are the battery current, terminal voltage, open-circuit voltage, and volume, respectively. Here, the trajectories of $I$ and $V$ depend on the battery usage. We select the worldwide harmonised light vehicle test procedure (WLTP) [32], which represents a wide range of vehicle driving conditions on urban, suburban, and highway roads, and scale it so that the maximum current becomes 1 C-rate for heat generation. The modified WLTP (mWLTP) heat generation profile is shown in Fig. 2(a). $V_{\text{ocv}}$ depends on the state of charge (SoC), and the relationship between $V_{\text{ocv}}$ and SoC has been determined using a lookup table for a typical LFP cell in [33].
Depending on where the cooling action is imposed, five different cooling scenarios are investigated, including surface cooling (SC), bottom tab cooling (bTC), bottom tab and surface cooling (bTSC), bottom and top tabs cooling (btTC), and all-tabs and surface cooling (aTSC). aTSC has all sides of the battery cell exposed to cooling and is representative of immersion cooling [34] where the cell is immersed in a dielectric fluid. Forced convection liquid cooling (for instance, using a cooling plate) via water or glycol is considered to apply on areas where tab or surface cooling occurs. This typically has a large convection coefficient $h$, which is set to be 400 Wm-2K-1, according to [4]. For areas not exposed to active cooling, mild air convection occurs, resulting in a small convection coefficient specified as $h$ = 30 Wm-2K-1.
All the simulations are conducted on a MacBook Pro with a 2.6GHz 6-Core Intel Core i7 processor and 16GB RAM. The first type of benchmark model is solved by the finite element method with an extremely fine mesh of 3610 triangular elements and is implemented in COMSOL Multiphysics v6.0 [35]. Our developed CSG models and their TEC alternatives are implemented in MATLAB R2023b. For simplicity in simulating the CSG models, the numbers of basis functions in the radial and axial directions, as introduced in (15), are chosen to be equal, i.e., $M=N$, and the model order $O=N^{2}$.
@@ IV-C Results of validation against the FEM model
The examination of the developed modelling method and their underlying assumptions for control-oriented applications is conducted for the battery cell and cooling scenarios specified in Section IV-B. Different CSG model orders, including $O$ = 1, 4, 9, 16, and 25, were evaluated. The results for cylindrical cell modelling compared with the ground truth represented by the FEM model are presented in Fig. 2.
Figure 2: Validation of the developed CSG thermal model for cylindrical battery cells. (a) Heat generation profile of the mWLTP. (b) Comparison of the CSG model of orders $O$ = 1 and 4, against the FEM model with the element number $N_{e}$ = 3610 and TEC model, under surface cooling (SC). Comparisons are done for the surface and core mid-point temperatures, i.e., $\hat{T}(1,0,t)$ and $\hat{T}(-1,0,t)$. (c) Comparison of the CSG thermal model against FEM and TEC models under the scenario of all-tabs and surface cooling (aTSC). (d) Maximum error of the CSG model with different orders under all cooling cases.
According to Fig. 2(d), the developed CSG model with an order of 25 can almost perfectly capture the battery cell’s distributed thermal dynamics under any of the considered cooling scenarios. The resulting modelling error is bounded by 0.03 C. Reducing the model order from 25 to 9, the maximum temperature deviation between the model’s output and its ground truth is 0.4 C within the entire test process. Similar observations are obtained for the studied pouch cell, in which the prediction error for the model with an order of 9 is less than 0.6 C. Remarkably, this high level of model fidelity is achieved in the presence of aggressive driving and cooling. The prediction error gradually increases when the model order is further lowered. This is consistent with Remark 2 and the applied finite-sum approximations in (14) and (23). Indeed, the model with an order of 1 or 4 can still well follow the trajectories of the true surface and core temperatures under different cooling scenarios, as seen in Fig. 2(b) and (c). Specifically, for SC, which involves one-side cooling, $O$ = 1 results in a maximum error of 1.26 C. For aTSC, which involves resolution in multiple directions (multi-side cooling), the same order produces a maximum error of 2.10 C.
The impact of current rates on model accuracy is also evaluated. To do so, the mWLTP profile is scaled by factors of two and three, resulting in maximum C-rates of 2C and 3C, respectively. Under these highly aggressive operating conditions and the SC scenario, the errors of the CSG model with different orders are summarized in Table I. The results demonstrate a clear necessity to increase the model order for accurately capturing battery cell thermal dynamics at elevated current rates due to pronounced thermal gradients and intensified spatial interactions. Considering a maximum permissible error of 1.5 C, model orders $O$ = 1, 9, and 16 are sufficient to faithfully model the thermal behaviour in the mWLTP, mWLTP×$\times$2, and mWLTP×$\times$3 scenarios, respectively. It is worth mentioning that with this relatively small magnitude of local errors obtained under all considered cooling scenarios and extreme battery usage conditions, the CSG model proves suitable for a wide range of battery management applications.
TABLE I: Effect of current rates on model accuracy.
Current profile $O$=1 $O$=4 $O$=9 $O$=16 $O$=25
mWLTP 1.26 0.46 0.13 0.09 0.03
mWLTP×$\times$2 3.83 2.36 1.11 0.98 0.73
mWLTP×$\times$3 4.89 3.62 1.85 1.36 1.01
@@
In addition to the above validation against the high-fidelity FEM model implemented in COMSOL, our future work will explore direct experiments to evaluate our model under various cooling configurations.